freshavacado850:

pl help im srugglin with this math question ss is down below im in 7th grade btw

justus:

Post your question, thanks.

freshavacado850:

Cookiedough1:

@jhonyy9

Cookiedough1:

He told me to tag him there

freshavacado850:

ok

freshavacado850:

@Ultrilliam

jhonyy9:

\[\frac{ 24x^2 +25x -47 }{ ax -2 } = -8x -3 - \frac{ 53 }{ ax -2 }\]

jhonyy9:

x not equal 2/a ,a is a constant

jhonyy9:

a = ?

jhonyy9:

so like a first step what will be the common denominator ?

freshavacado850:

wuts a comman dominato?

jhonyy9:

or denominator common

freshavacado850:

wuts demomnitae comman?

jhonyy9:

somebody know ?

freshavacado850:

idk im a dumbo

jhonyy9:

on the left and on the right side there are two fractions with denominator the same so in this way the common denominator will be ax -2

jhonyy9:

so now what we need to do with these terms without denominators ?

jhonyy9:

for example how you add two fractions 1/x + 3/2 = bc. the denominator common is 2x the numerator of the first fraction we need multiplie by 2 and the numerator of second fraction by x and will get = (2 +3x)/2x

freshavacado850:

ok bet thanks

freshavacado850:

you smart boi

jhonyy9:

so now using this example what we need to do with these terms without denominator ?

freshavacado850:

oh ok thats ok i dont need help manymore

freshavacado850:

thanks for your help